OpenStudy (mb194227):

How do I turn a quadratic equation in standard form into a quadratic equation in vertex form?

OpenStudy (mathmale):

Let's begin with standard form: y=ax^2 + bx + c Our goal is to end up with an equation of this form: y=a(x-h)^2+k, where (h,k) is the vertex. Consider how you would find the x- and y-coordinates of the vertex if given y = ax^2 + bx + c. Hint: use part (not all) of the quadratic formula: x=-b/(2a)

OpenStudy (mathmale):

This x would be your h in (h,k). Use the original equation to find y when x=h.